OpenStudy (anonymous):

consider the line whose equation is 4x-8y=8 A.) write the equation in slope-intercept form B.) write a slope of a line perpendicular to this line C.) write the equation of the line, in point-slope form, that is perpendicular to the line that passes through the point (1, 2)

OpenStudy (anonymous):

@texaschic101 @Destinymasha @Goku-Kai

OpenStudy (anonymous):

you have to write an answer for each

OpenStudy (destinymasha):

shoot -- im not the best at math lol

OpenStudy (anonymous):

thats okay I was just wondering if maybe you could help

OpenStudy (anonymous):

a) y=1/2x-1 b) y=-2/1-1

OpenStudy (anonymous):

Refer to the Mathematica 9 Home Edition solution attachment which includes a plot.